OpenStudy (anonymous):

what are the steps to solve this problem 0.000027÷0.000009 and the answer in simplest form

OpenStudy (loser66):

0.000027 = 27 *10^(-6) 0.000009 = 9*10^(-6) -------------------------- = 3